# Does DetectGPT Fully Utilize Perturbation? Bridge Selective Perturbation to Fine-tuned Contrastive Learning Detector would be Better

### v3

@@ 4.5.2 Analysis on Selective Strategies
Our method is a neat perturbation strategy that is mainly for the MGT detection task and empirically shows great performance. Beyond the Pecola and random perturbation method, we experiment with two other perturbation strategies: rank-based perturbation and keyword-based perturbation. In rank-based perturbation, we use the rescaled rank of next-token probability on GPT2-medium as the weight for perturbation position selection. In keyword-based perturbation, we prevent changes in the keywords extracted by the VLT-5 language model Pęzik et al. (2022) during perturbation. As shown in Table 5, the experimental results of selective perturbation outperform the random perturbation method by 1.20%, 2.04%, and 2.49% in average accuracy on the 64-example GPT2 dataset. The Importance-based strategy of Pecola is the highest.
@@
We further find that the random strategies of DetectGPT lead to more masking-filling failures than selective ones, which cause execution errors. Mask-filling failure ratio in Table 6 indicates that selective strategy based on token importance greatly decreases the failure ratio by 3.64%.
@@ 4.5.3 Generalization on Mask-Filling Models
We study the influence of various mask-filling models on the performance of Pecola, including Bert (110M; Devlin et al. 2019), Bart (139M; Mike et al. 2020), GPT-2 (380M; Radford et al. 2019), Twhin-bert (279M; Zhang et al. 2023), XLM (279M; Alexis et al. 2020), XLNet (110M; Yang et al. 2019), RoBERTa(135M; Liu et al. 2019), and LLaMa-2(7B; Touvron et al. 2023). As depicted in Fig. 3, the results of all mask-filling models, in accuracy, surpass the baseline. Furthermore, the fluctuation of Pecola’s performance across different mask-filling models is relatively slight. It confirms that Pecola is not reliant on a specific filling model, showing great generalization capability. The experiment details and results are in Appendix D.1.
@@ 4.5.4 Generalization on Data
Cross-domain. We evaluate Pecola on the HC3 dataset, the meta-information details are in Appendix A.2. For the three domains of data, we use one of them as training data (64-shot), and the remaining domains of data as testing data. The results in Table 7 show that Pecola is more effective than the best baseline and SOTA method on average. For example, compared to Roberta, Pecola outperforms by 4.61% in three domains on average. And Pecola maintains a 1.63% higher accuracy on average than SOTA DetectGPT.
| Domain | Medicine | Finance | Computer Sci. | Average |
@@
Cross-genre. We generalize Pecola between News articles (GPT3.5 dataset) and QA answers (HC3 dataset) on the 64-shot settings. As shown in Table 8, when the GPT-3.5 dataset is the training set, Pecola outperforms by 10.21%; and when the HC3 dataset is the training set, Pecola outperforms by 6.98% to the best competitor.
@@ 5 Conclusion
In this paper, we introduce Pecola, a novel machine-generated text detection method that effectively bridges and integrates metric-based and fine-tuned detectors for MGT detection. To relieve the information loss caused by the random masking used in DetectGPT, we present a token-level selective strategy perturbation method. To more fully distinguish meaningful recombination spaces and reduce reliance on the mask-filling models, we present a token-level weighted multi-pairwise contrastive learning method. In few-shot settings, experimental results show that Pecola significantly enhances the performance of PLMs in MGT detection. Subsequent analytical experiments validate Pecola’s effectiveness, robustness, generalization, and capability in detecting short texts.
@@ A.1 Hyperparameter Details
A.2 Meta-information
We evaluate Pecola effectiveness from domains and genres on the HC3 dataset, which primarily includes Medicine, Finance, and Computer Science domain QA, as shown in Table 10.
| Domain | Medicine | Finance | Computer Science |
@@ B.2 Score Threshold
Appendix C Efficiency and Scalability of Pecola
@@
Appendix D Further Clarification on Perturbation
Conversely, diversity assesses the range and variability of perturbed data, utilizing metrics Dist-1 and Dist-2 Celikyilmaz et al. (2020). Here, we use three common perturbation methods to demonstrate the importance of not arbitrarily changing important tokens and the significance of select masks. (1) Token Substitution Zhang et al. (2015), replaces tokens with synonyms from WordNet Miller (1992); (2) SwitchOut Wang et al. (2018), uniformly samples and randomly substitutes from the vocabulary of test samples; and (3) Two-stage Wei et al. (2021) trains the mask-filling model on the original data.
The ideal perturbation result is to have high Affinity scores while ensuring high Diversity scores Celikyilmaz et al. (2020). As shown in Table 12, through Selective Strategy Perturbation, models achieve better diversity with high distribution shifts. And the overall improvement in Affinity by over 20% also shows greater diversity than the original data. The above results demonstrate the superiority of our perturbation method.
| Method | Affinity | Diversity | |
| Dist-1 | Dist-2 | ||
| Original | - | 8.70 | 50.32 |
| Token Substitution | -20.00 | 3.38 | 43.43 |
| SwitchOut | -22.06 | 6.81 | 53.61 |
| Two-stage | -21.13 | 3.24 | 41.85 |
| Pecola Mask-Filling | -1.34 | 15.59 | 57.01 |
D.1 Impact of the Chosen Mask-filling Models
@@
Appendix E Detecting Shorter Texts

### v4

@@ 4.5.2 Analysis on Selective Strategies
In this section, we compare various strategies for selection in Pecola. Beyond the Pecola’s importance-based perturbation method and random perturbation method (DetectGPT), we experiment with two other perturbation strategies: rank-based perturbation and keyword-based perturbation. In rank-based perturbation, we use the rescaled rank of next-token probability on GPT2-medium as the weight for perturbation position selection. In keyword-based perturbation, we prevent changes in the keywords extracted by the VLT-5 model Pęzik et al. (2022) during perturbation. As shown in Table 6, the experimental results of selective perturbation outperform the random perturbation method by 1.20%, 2.04%, and 2.49% in average accuracy on the 64-example GPT2 dataset. And the importance-based strategy is the highest.
@@
Further, we test the mask-filling failure ratio across the above strategies to interpret our outperformance. We find that the random strategy leads to more masking-filling failures than selective ones, which cause execution errors. Results in Table 7 indicate that selective strategy based on token importance performs the best, decreasing the failure ratio by 3.64% than random.
@@ 4.5.3 Generalization on Mask-Filling Models
We study the influence of various mask-filling models on the performance of Pecola, including Bert (110M; Devlin et al. 2019), Bart (139M; Mike et al. 2020), GPT-2 (380M; Radford et al. 2019), Twhin-bert (279M; Zhang et al. 2023), XLM (279M; Alexis et al. 2020), XLNet (110M; Yang et al. 2019), RoBERTa (125M; Liu et al. 2019), and LLaMA-2 (7B; Touvron et al. 2023). As depicted in Fig. 3, the results of all mask-filling models surpass the baseline in terms of accuracy. Furthermore, the fluctuation of Pecola’s performance across different mask-filling models is relatively slight. It confirms that Pecola is not reliant on a specific filling model, showing great generalization capability. The remaining full experimental results of different mask-filling models are in Appendix C.2.
@@ 4.5.4 Generalization on Data
Cross-domain. We evaluate Pecola on the HC3 dataset crossing three QA domains, namely Medicine, Finance, and Computer Science. The meta-information details are in Appendix A.2. For the three domains of data, we use one of them as training data (64-shot), and the remaining domains of data as testing data. The results in Table 8 show that Pecola is more effective than the best baseline and SOTA method on average. For example, compared to Roberta, Pecola outperforms by 4.61% in three domains on average. And Pecola maintains a 1.63% higher accuracy on average than SOTA DetectGPT.
| Domain | Medicine | Finance | Comp. Sci. | Average |
@@
Cross-generator. We generalize Pecola between News articles (GPT3.5 dataset) and QA answers (HC3 dataset) on the 64-shot settings. As shown in Table 9, when the GPT-3.5 dataset is the training set, Pecola outperforms by 10.21%; and when the HC3 dataset is the training set, Pecola outperforms by 6.98% to the best competitor.
@@ 5 Conclusion
In this paper, we introduce Pecola, a novel machine-generated text detection method that effectively bridges and integrates metric-based and fine-tuned detectors for MGT detection. To relieve the information loss caused by the random masking used in DetectGPT, we present a token-level selective strategy perturbation method. To better distinguish meaningful recombination spaces and reduce reliance on the mask-filling models, we present a token-level weighted multi-pairwise contrastive learning method. In few-shot settings, experimental results show that Pecola significantly enhances the performance of PLMs in MGT detection. Subsequent analytical experiments validate Pecola’s effectiveness, robustness, generalization, and capability in detecting short texts.
Acknowledgements
We thank all the anonymous reviewers and the area chair for their helpful feedback, which aided us in greatly improving the paper. This work is supported by National Key R&D Program (2023YFB3107400), National Natural Science Foundation of China (62272371, 62103323, U21B2018, 62161160337, U20B2049), Initiative Postdocs Supporting Program (BX20190275, BX20200270), China Postdoctoral Science Foundation (2019M663723, 2021M692565), Fundamental Research Funds for the Central Universities under grant (xzy012024144), and Shaanxi Province Key Industry Innovation Program (2021ZDLGY01-02).
@@ A.1 Hyperparameter Details
A.2 Dataset Meta Information
We evaluate Pecola effectiveness from domains and generators on the HC3 dataset, which primarily includes Medicine, Finance, and Computer Science domain QA, as shown in Table 11.
| Domain | Medicine | Finance | Comp. Sci. |
@@ B.2 Score Threshold
Appendix C Efficiency of Pecola
C.1 Scalability of Base Models at Different Scales
@@
C.2 Impact of the Chosen Mask-filling Models
@@
| Dataset | Shot | Metric | RoBERTa | GLTR | CE+SCL | CE+Margin | IT:Clust | CoCo | DetectGPT | Fast-Detect. | Pecola |
| Grover | 10000 | Acc | 86.13 | 60.40 | 86.57 | 86.25 | 72.65 | 85.23 | 61.42 | 65.49 | 86.70 |
| F1 | 84.07 | 59.82 | 84.95 | 85.10 | 63.21 | 83.67 | 54.28 | 63.29 | 86.66 | ||
| GPT-2 | 10000 | Acc | 89.56 | 77.55 | 90.19 | 90.30 | 81.65 | 89.78 | 78.74 | 80.06 | 91.10 |
| F1 | 89.51 | 76.39 | 90.15 | 90.27 | 81.54 | 89.01 | 71.13 | 80.64 | 91.10 | ||
| GPT-3.5 | 7000 | Acc | 99.89 | 93.50 | 99.74 | 99.90 | 99.09 | 99.44 | 90.80 | 94.72 | 99.95 |
| F1 | 99.89 | 93.58 | 99.74 | 99.90 | 99.09 | 99.44 | 89.14 | 94.76 | 99.95 | ||
| HC3 | 10000 | Acc | 99.84 | 98.39 | 99.89 | 99.86 | 98.80 | 99.46 | 95.13 | 98.32 | 99.92 |
| F1 | 99.84 | 98.49 | 99.89 | 99.86 | 98.80 | 99.46 | 95.05 | 98.02 | 99.92 |
C.3 Further Experiments on Full Datasets
To demonstrate Pecola’s superiority over the whole training set, we conduct a more in-depth test, as shown in Table 14. We train the detector using 10,000 samples from the Grover, GPT-2, and HC3 datasets, and 7,000 samples from GPT-3.5 as our training sets. Comparatively, Pecola outperforms the second-best results in accuracy and F1-score by 0.13% and 1.56%, 0.80% and 0.83%, 0.05% and 0.05%, 0.03% and 0.03% respectively, across four datasets.